# Avoiding Catastrophe in Online Learning by Asking for Help

### v3

@@ Abstract
Most learning algorithms with formal regret guarantees assume that no mistake is irreparable and essentially rely on trying all possible behaviors. This approach is problematic when some mistakes are catastrophic, i.e., irreparable. We propose an online learning problem where the goal is to minimize the chance of catastrophe. Specifically, we assume that the payoff in each round represents the chance of avoiding catastrophe that round and aim to maximize the product of payoffs (the overall chance of avoiding catastrophe) while allowing a limited number of queries to a mentor. We first show that in general, any algorithm either constantly queries the mentor or is nearly guaranteed to cause catastrophe. However, in settings where the mentor policy class is learnable in the standard online learning model, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows. Conceptually, if a policy class is learnable in the absence of catastrophic risk, it is learnable in the presence of catastrophic risk if the agent can ask for help.
@@ 1 Introduction
There has been mounting concern over catastrophic risk from AI, including but not limited to autonomous weapon accidents [1], bioterrorism [30], and cyberattacks on critical infrastructure [14]. See [10] and [20] for taxonomies of societal-scale AI risks. In this paper, we use “catastrophe” to refer to any kind of irreparable harm. This definition also covers smaller-scale (yet still unacceptable) incidents such as serious medical errors [11], crashing a robotic vehicle [24], or discriminatory sentencing [39].
@@
More precisely, trial-and-error-style algorithms only work when catastrophe is assumed to be impossible. This assumption manifests differently in different subtypes of online learning. In the standard online learning model, the agent’s actions have no permanent effect on the environment. [^1] More precisely, the state can depend on the agent’s previous actions, but the agent’s performance is always evaluated with respect to the optimal policy on the same sequence of states. Online reinforcement learning allows the agent’s actions to permanently affect the environment, but typically assumes that either no action has irreversible effects (e.g., [21]) or that the agent is reset at the start of each “episode” (e.g., [2]). One could train an agent entirely in a controlled lab setting where the above assumptions do hold, but we argue that sufficiently general agents will inevitably encounter novel scenarios when deployed in the real world. Machine learning models often behave unpredictably in unfamiliar environments (see, e.g., [32]), and we do not want AI biologists or robotic vehicles to behave unpredictably.
The goal of this paper is to understand the conditions under which it is possible to formally guarantee avoidance of catastrophe in online learning. Certainly some conditions are necessary, because if the agent can only learn by trying actions directly, the problem is hopeless: any untried action could lead to paradise or disaster and the agent has no way to predict which. In the real world, however, one needn’t learn through pure trial-and-error: one can also ask for help.
@@ 1.1 Our model
We propose an online learning model of avoiding catastrophe with mentor help. On each time step, the agent observes a state, selects an action (or queries the mentor), and obtains a payoff. [^2] Depending on the field, states are also known as inputs, instances, or contexts, actions are also known as outputs, labels, or arms, and payoffs are also known as rewards or losses. Each payoff represents the probability of avoiding catastrophe on that time step, and the agent’s goal is to maximize the product of payoffs, which corresponds to minimizing the total chance of catastrophe. We consider the product of payoffs obtained while learning, not the product of payoffs of some final policy: finding a catastrophe-avoiding policy is futile if catastrophe occurs while learning.
The (possibly suboptimal) mentor has a fixed policy, and when queried, the mentor illustrates their policy’s action in the current state. We desire an agent whose regret – defined as the gap between the mentor’s performance and the agent’s performance – approaches zero as the time horizon $T$ grows. In other words, with enough time, the agent should avoid catastrophe nearly as well as the mentor. We also expect the agent to become self-sufficient over time: formally, the number of queries to the mentor should be sublinear in $T$, or equivalently, the rate of querying the mentor should go to zero.
@@ 1.2 Our assumptions
The agent needs some way to make inferences about unseen states in order to decide when to ask for help. Much past work has used Bayesian inference, which suffers tractability issues in complex environments. [^3] For the curious reader, [4] provides a thorough treatment. See also Section 2. In this paper, we instead assume that the mentor policy satisfies a Lipschitz-flavored condition we call “local generalization”: informally, if the mentor told us that action $a$ was safe in a nearby state, then $a$ is probably (but not certainly) also safe in the current state. Local generalization captures the intuition that one can transfer knowledge between similar states but the generalization becomes progressively less reliable as the states become less similar. Unlike Bayesian inference, local generalization only requires computing distances and is compatible with any state space which admits a distance metric.
Unlike the standard online learning model, we assume that the agent does not observe payoffs. This is because the payoff in our model represents the chance of avoiding catastrophe on that time step. In the real world, one only observes whether catastrophe occurred, not its probability. [^4] One may be able to detect “close calls” in some cases, but observing the precise probability seems unrealistic.
@@ 1.3 Standard online learning
An overview of standard online learning is in order before discussing our results. In the standard model, the agent observes a state on each time step and must choose an action. An adversary then reveals the correct action, which results in some payoff for the agent. The goal is sublinear regret, i.e., the average regret per time step should go to 0 as $T\to\infty$. Figure 1 delineates the precise differences between the standard model and our model.
If the adversary’s choices are unconstrained, the problem is hopeless: if the adversary determines the correct action on each time step randomly and independently, the agent can do no better than random guessing. However, sublinear regret becomes possible if (1) the hypothesis class has finite Littlestone dimension [26], or (2) the hypothesis class has finite VC dimension [38] and the input is $\sigma$-smooth [^5] Informally, the adversarial chooses a distribution over states instead of a precise state. See Section 3. [17].
@@
Objective Regret goal Feedback Mentor Local gen.
@@
Figure 1: Comparison between the standard online learning model and our model. On a technical level, sublinear vs subconstant regret is a more important distinction than the sum vs product of payoffs: for example, taking a logarithm can transform the product into a sum. However, the multiplicative objective and the goal of subconstant regret are tightly related, since the former’s interpretation as the chance of catastrophe is what gives rise to the latter: we want the total (not average) chance of catastrophe to go to 0.
@@ 1.4 Our results
Informally, we show that avoiding catastrophe with the help of a mentor and local generalization is no harder than online learning without catastrophic risk.
More precisely, we first show that in general, any algorithm with sublinear queries to the mentor has arbitrarily poor regret in the worst-case (Theorem 4.1). As a corollary, even when it is possible to avoid catastrophe with certainty, any algorithm either asks for help on a constant fraction of time steps or is nearly guaranteed to cause catastrophe. Unlike online learning where the general impossibility result is trivial (the agent might as well guess randomly given an unconstrained adversary), local generalization significantly limits the adversary’s power and necessitates a careful construction.
Next, we present a simple algorithm whose total regret and rate of querying the mentor both go to 0 as $T\to\infty$ when either (1) the mentor policy class has finite Littlestone dimension or (2) the mentor policy class has finite VC dimension and the state sequence is $\sigma$-smooth. Our algorithm can handle a multi-dimensional unbounded state space and does not need detailed access to the feature embedding, instead using two simple operations. It does need to know the mentor policy class, as is standard in online learning. We initially prove the theorem for binary actions (Theorem 5.2) and then reduce learning with (finitely) many actions to the binary action case (Theorem C.1).
Finally, Appendix D presents another algorithm with the same guarantees of subconstant regret and sublinear queries under different assumptions. Most importantly, this result applies to the class of thresholds on $[0,1]$, which is known to be hard for standard online learning (see Section 2). This shows that our problem and standard online learning do not have exactly the same difficulty, but rather our problem is no harder than online learning.
The rest of the paper is structured as follows. Section 2 discusses related work. Section 3 formally defines our model. Section 4 presents our negative result for general mentor policies. Section 5 presents our positive result for simple mentor policy classes. Proofs are deferred to the appendix.
@@ 2 Related work
Despite the ubiquity of irreparable/irreversible costs in the real world, theoretical work on this topic remains limited. This may be due to the fundamental modeling question of how the agent should learn about novel states or actions without actually trying them.
The most common approach is to allow the agent to ask for help. This alone is insufficient, however: the agent must have some way to decide when to ask for help. A popular solution is to perform Bayesian inference on the world model, but this has two tricky requirements: (1) a prior distribution which contains the true world model (or an approximation), and (2) an environment where computing (or approximating) the posterior is tractable. A finite set of possible environments satisfies both conditions, but is unrealistic in many real-world scenarios. In contrast, our algorithm can handle an uncountable policy class and a continuous unbounded state space, which is crucial for many real-world scenarios in which one never sees the exact same state twice.
Bayesian inference combined with asking for help is studied by [8; 9; 25; 28]. We also mention [15; 29; 37], which utilize Bayesian inference in the context of safe (online) reinforcement learning without asking for help (and without regret bounds).
We are only aware of two papers which theoretically address irreversibility without Bayesian inference: [13] and [27]. The former proposes to sample trajectories and learn reversibility based on temporal consistency between states: intuitively, if $s_{1}$ always precedes $s_{2}$, we can infer that $s_{1}$ is unreachable from $s_{2}$. Although the paper theoretically grounds this intuition, there is no formal regret guarantee. The latter presents an algorithm which asks for help in the form of rollouts from the current state. However, the regret bound and number of rollouts are both linear in the worst case, due to the dependence on the $\gamma^{*}$ parameter which roughly captures how bad an irreversible action can be. In contrast, our algorithm achieves good regret even when actions are maximally bad.
To our knowledge, we are the first to provide an algorithm which formally guarantees avoidance of catastrophe (with high probability) without Bayesian inference.
@@
See [7] and Chapter 21 of [34] for introductions to online learning. A classical result states that sublinear regret is possible iff the hypothesis class has finite Littlestone dimension [26]. However, even some simple hypothesis classes have infinite Littlestone dimension, such as the class of thresholds on $[0,1]$ (Example 21.4 in [34]). Recently, [17] showed that if the adversary only chooses a distribution over states rather than the precise state, only the weaker assumption of finite VC dimension [38] is needed for sublinear regret. Specifically, they assume that each state is sampled from a distribution whose concentration is upper bounded by $\frac{1}{\sigma}$ times the uniform distribution. This framework is known as smoothed analysis.
Smoothed analysis.
Smoothed analysis was originally proposed by [36] to explain the fact that some algorithms almost always perform well in practice, despite the theoretical existence of pathological inputs. Informally, algorithms which perform well under smoothed analysis are guaranteed to perform well for nearly all – but not necessarily all – inputs.
Active learning.
Our assumption that the agent only receives feedback in response to queries falls under the umbrella of active learning [18]. This contrasts with passive learning, where the agent receives feedback automatically. Although ideas from active learning could be useful in our domain, we are not aware of any results from that literature which account for irreversible costs.
@@
Perhaps the best-studied example of a multiplicative objective is the Nash welfare function (also known as the Nash bargaining solution), defined as the geometric mean of agent utilities [23; 31]. Inspired by Nash welfare, [3] recently studied algorithms for maximizing the geometric mean of payoffs across time steps in the multi-armed bandit problem. However, the geometric mean of payoffs is much more forgiving of errors compared to the product of payoffs. Imagine that the agent obtains a payoff of $1-1/\sqrt{T}$ per time step: then $\lim_{T\to\infty}(1-1/\sqrt{T})^{T}=0$ but $\lim_{T\to\infty}((1-1/\sqrt{T})^{T})^{1/T}=\lim_{T\to\infty}(1-1/\sqrt{T})=1$. In our setting, it is crucial that the total chance of catastrophe (the product of payoffs) is small, not just the average chance of catastrophe per time step (the geometric mean of payoffs). This is closely related to the difference between sublinear and subconstant regret discussed in Section 1.3.
@@ 3 Model
States.
Let $\mathbb{N}$ refer to the set of strictly positive integers and let $T\in\mathbb{N}$ be the time horizon. Let $S\subseteq\mathbb{R}^{n}$ be the state space and let $\mathbf{s}=(s_{1},s_{2},\dots s_{T})\in S^{T}$ be the sequence of states. In the adversarial setting, each $s_{t}$ can have arbitrary dependence on the events of prior time steps. In the smoothed setting, the adversary only chooses the distribution $s_{t}$ from which $s_{t}$ is sampled. Formally, a distribution $\mathcal{D}$ over $S$ is $\sigma$-smooth if for any $X\subseteq S$, $\mathcal{D}(X)\leq\frac{1}{\sigma}U(X)$. (In the smoothed setting, we assume that $S$ supports a uniform distribution $U$ [^6] For example, $S$ having finite Lebesgue measure is sufficient. Note that this does not imply boundedness. Alternatively, $\sigma$-smoothness can be defined with respect to a different distribution, as long as the Radon-Nikodym derivative is uniformly bounded; see Definition 1 of [5].) If each $s_{t}$ is sampled from a $\sigma$-smooth $\mathcal{D}_{t}$, we say that $\mathbf{s}$ is $\sigma$-smooth. The sequence $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ can still be adaptive, i.e., the choice of $\mathcal{D}_{t}$ can depend on the events of prior time steps.
Actions.
Let $A$ be a finite set of actions. There also exists a special action $\hat{a}$ which corresponds to querying the mentor. For $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$. On each time step $t\in[T]$, the agent must select an action $a_{t}\in\ A\cup\{\hat{a}\}$, which generates payoff $\mu(s_{t},a_{t})\in[0,1]$.
@@
The mentor is endowed with a (not necessarily optimal) policy $\pi^{m}:S\to A$. When action $\hat{a}$ is chosen, the mentor informs the agent of the action $\pi^{m}(s_{t})$ and the agent obtains payoff $\mu(s_{t},\pi^{m}(s_{t}))$. For brevity, let $\mu^{m}(s)=\mu(s,\pi^{m}(s))$. The agent never observes payoffs: the only way to learn about $\mu$ is by querying the mentor.
We would like an algorithm which becomes “self-sufficient” over time: the rate of querying the mentor should go to 0 as $T\to\infty$. Intuitively, this is equivalent to the cumulative number of queries being sublinear in $T$: $|Q_{T}|\in o(T)$, where $Q_{T}=\{t\in[T]:a_{t}=\hat{a}\}$. However, interpreted literally, $|Q_{T}|\in o(T)$ is not a rigorous statement. This is because $|Q_{T}|$ is not just a function of $T$: it can also depend on $\mathbf{s}$, $\mu$, $\pi^{m}$, and any randomness in the agent’s actions.
To formalize this, let $Q_{T}(\mathbf{s},\mu,\pi^{m})$ be the set of time steps with queries given $\mathbf{s},\mu$, and $\pi^{m}$. Then we say that $|Q_{T}|$ is sublinear in $T$, denoted $|Q_{T}|\in o(T)$, if there exists $g:\mathbb{N}\to\mathbb{N}$ such that $g(T)\in o(T)$ and $|Q_{T}(\mathbf{s},\mu,\pi^{m})|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{m}$ (and for all possible realizations of the agent’s randomization). However, we will mostly write $Q_{T}$ and $|Q_{T}|\in o(T)$ for brevity.
@@
We assume that the mentor policy permits “local generalization”. Formally, we assume there exists $L>0$ such that for all $s,s^{\prime}\in S$, $|\mu^{m}(s)-\mu(s,\pi^{m}(s^{\prime}))|\leq L||s-s^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance. Informally, if the agent is in state $s$, taking the mentor action from a nearby state $s^{\prime}$ is almost as good as taking the mentor action for $s$. In other words, it’s okay to use an action you learned from a nearby state.
This assumption might initially feel less intuitive than a standard Lipschitz continuity assumption of $|\mu(s,a)-\mu(s^{\prime},a)|\leq L||s-s^{\prime}||$. However, Lipschitz continuity is strictly stronger than local generalization when the mentor is optimal (Proposition E.1). The properties are incomparable in general, but Lipschitz continuity allows odd situations where the mentor provides an action which is only acceptable in that specific state. This sets the agent up for failure: even if such actions exist, a thoughtful mentor should avoid those actions when possible. See Figure 2 for an illustration.
[…]
Figure 2: The difference between local generalization and Lipschitz continuity. The x-axis corresponds to a one-dimensional state space, with $s_{1}$ and $s_{2}$ denoting two nearby states. The y-axis is the payoff. In all figures, $\pi^{m}(s_{1})=0$ and $\pi^{m}(s_{2})=1$. Left: What one might consider the “typical case” with a near-optimal mentor. The mentor takes action 0 in $s_{1}$, so $|\mu(s_{2},0)-\mu(s_{2},\pi^{m}(s_{2})|$ is small (local generalization) and $|\mu(s_{2},0)-\mu(s_{1},0)|$ is small (Lipschitz continiuty). Middle: Local generalization without Lipschitz continuity. Action 0 is much worse in $s_{2}$, but it is still nearly as good as the mentor’s action in $s_{2}$, which is also quite bad. We primarily care about the agent’s performance relative to the mentor. Right: Lipschitz continuity without local generalization. The payoff of each action is similar in nearby states. However, despite the mentor recommending action 0 in $s_{1}$, action 0 is much worse than the mentor’s action in $s_{2}$. The mentor is setting up the agent for failure by recommending action 0 in $s_{1}$.

### v4

@@ Abstract
Most learning algorithms with formal regret guarantees assume that all mistakes are recoverable and essentially rely on trying all possible behaviors. This approach is problematic when some mistakes are catastrophic, i.e., irreparable. We propose an online learning problem where the goal is to minimize the chance of catastrophe. Specifically, we assume that the payoff in each round represents the chance of avoiding catastrophe that round and try to maximize the product of payoffs (the overall chance of avoiding catastrophe) while allowing a limited number of queries to a mentor. We first show that in general, any algorithm either constantly queries the mentor or is nearly guaranteed to cause catastrophe. However, in settings where the mentor policy class is learnable in the standard online model, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows. Conceptually, if a policy class is learnable in the absence of catastrophic risk, it is learnable in the presence of catastrophic risk if the agent can ask for help.
@@ 1 Introduction
There has been mounting concern over catastrophic risk from AI, including but not limited to autonomous weapon accidents (Abaimov & Martellini, 2020), bioterrorism (Mouton et al., 2024), and cyberattacks on critical infrastructure (Guembe et al., 2022). See Critch & Russell (2023) and Hendrycks et al. (2023) for taxonomies of societal-scale AI risks. In this paper, we use “catastrophe” to refer to any kind of irreparable harm. This definition also covers smaller-scale (yet still unacceptable) incidents such as serious medical errors (Rajpurkar et al., 2022), crashing a robotic vehicle (Kohli & Chadha, 2020), and discriminatory sentencing (Villasenor & Foggo, 2020).
@@
More precisely, trial-and-error-style algorithms only work when catastrophe is assumed to be impossible. This assumption manifests differently in different subtypes of online learning. In the standard online learning model, the agent’s actions have no permanent effect on the environment. [^1] More precisely, the input can depend on the agent’s previous actions, but the agent’s performance is always evaluated relative to the optimal policy’s performance on the same sequence of inputs. Online reinforcement learning allows the agent’s actions to permanently affect the environment, but typically assumes that either no action has irreversible effects (e.g., Jaksch et al. (2010)) or that the agent is reset at the start of each “episode” (e.g., Azar et al. (2017)). One could train an agent entirely in a controlled lab setting where the above assumptions do hold, but we argue that sufficiently general agents will inevitably encounter novel scenarios when deployed in the real world. Machine learning models often behave unpredictably in unfamiliar environments (see, e.g., Quionero-Candela et al. (2009)), and we do not want AI biologists or robotic vehicles to behave unpredictably.
The goal of this paper is to understand the conditions under which it is possible to formally guarantee avoidance of catastrophe in online learning. Certainly some conditions are necessary, because if the agent can only learn by trying actions directly, the problem is hopeless: any untried action could lead to paradise or disaster and the agent has no way to predict which. In the real world, however, one needn’t learn through pure trial-and-error: one can also ask for help. We think it is critical for high-stakes AI applications to employ a designated supervisor who can be asked for help. Examples include a human doctor supervising AI doctors, a robotic vehicle with a human driver who can take over in emergencies, autonomous weapons with a human operator, and many more. We hope that our work constitutes a small step in the direction of practical safety guarantees for such applications.
@@ 1.1 Our model
We propose an online learning model of avoiding catastrophe with mentor help. On each time step, the agent observes an input, selects an action (or queries the mentor), and obtains a payoff. Each payoff represents the probability of avoiding catastrophe on that time step (conditioned on no prior catastrophe). The agent’s goal is to maximize the product of payoffs, which is equal to the overall probability of avoiding catastrophe. [^2] Conditioning on no prior catastrophe means we do not need to assume that these probabilities are independent (and if catastrophe has already occurred, this time step does not matter). This is due to the chain rule of probability. As is standard in online learning, we consider the product of payoffs obtained while learning, not the product of payoffs of some final policy.
The (possibly suboptimal) mentor has a fixed policy, and when queried, the mentor illustrates their policy’s action for the current input. We desire an agent whose regret – defined as the gap between the mentor’s performance and the agent’s performance – approaches zero as the time horizon $T$ grows. In other words, with enough time, the agent should avoid catastrophe nearly as well as the mentor. We also expect the agent to become self-sufficient over time: formally, the number of queries to the mentor should be sublinear in $T$, or equivalently, the rate of querying the mentor should go to zero.
@@ 1.2 Our assumptions
The agent needs some way to make inferences about unqueried inputs in order to decide when to ask for help. Much past work has used Bayesian inference, which suffers tractability issues in complex environments. [^3] For the curious reader, Betancourt (2018) provides a thorough treatment. See also Section 2. We instead assume that the mentor policy satisfies what we call local generalization: informally, if the mentor told us that an action was safe for a similar input, then that action is probably also safe for the current input (see Section 3 for a formal definition and further discussion). This captures the intuition that one can transfer knowledge between similar situations. Unlike Bayesian inference, local generalization only requires computing distances and is compatible with any input space which admits a distance metric.
Unlike the standard online learning model, we assume that the agent does not observe payoffs. This is because the payoff in our model represents the chance of avoiding catastrophe on that time step. In the real world, one only observes whether catastrophe occurred, not its probability. [^4] One may be able to detect “close calls” in some cases, but observing the precise probability seems unrealistic. One consequence of this is that the payoff function does not affect the behavior of the algorithm, so our positive results hold simultaneously for all payoff functions which satisfy local generalization.
@@ 1.3 Standard online learning
An overview of standard online learning is in order before discussing our results. In the standard model, the agent observes an input on each time step and must choose an action. An adversary then reveals the correct action, which results in some payoff to the agent. The goal is sublinear regret with respect to the sum of payoffs, or equivalently, the average regret per time step should go to 0 as $T\to\infty$. Figure 1 delineates the precise differences between the standard model and our model.
If the adversary’s choices are unconstrained, the problem is hopeless: if the adversary determines the correct action on each time step randomly and independently, the agent can do no better than random guessing. However, sublinear regret becomes possible if (1) the hypothesis class has finite Littlestone dimension (Littlestone, 1988), or (2) the hypothesis class has finite VC dimension (Vapnik & Chervonenkis, 1971) and the input is $\sigma$-smooth [^5] Informally, the adversary chooses a distribution over inputs instead of a precise input. See Section 3. (Haghtalab et al., 2024).
@@
Objective Regret goal Feedback Mentor Local generalization
@@
Figure 1: Comparison between the standard online learning model and our model.
@@ 1.4 Our results
At a high level, we show that avoiding catastrophe with the help of a mentor and local generalization is no harder than online learning without catastrophic risk.
More precisely, we first show that in general, any algorithm with sublinear queries to the mentor has arbitrarily poor regret in the worst-case (Theorem 4.1). This means that even when the mentor can avoid catastrophe with certainty, any algorithm either needs excessive supervision or is nearly guaranteed to cause catastrophe. Unlike online learning where the general impossibility result is trivial (the agent might as well guess randomly given an unconstrained adversary), local generalization significantly limits the adversary’s power and necessitates a careful analysis.
Next, we present a simple algorithm whose total regret and rate of querying the mentor both go to 0 as $T\to\infty$ when either (1) the mentor policy class has finite Littlestone dimension or (2) the mentor policy class has finite VC dimension and the input sequence is $\sigma$-smooth. Our algorithm can handle a unbounded input space and does not need detailed access to the feature embedding, instead relying on two simple operations. It does need to know the mentor policy class, as is standard in online learning. We initially prove the theorem for binary actions (Theorem 5.2) and then reduce learning with many actions to the binary action case (Theorem C.1).
Along the way, we prove that the same subconstant bound holds for standard additive regret (Theorem 5.3). Essentially, our techniques are equally effective for maximizing the sum of payoffs and the product of payoffs. We emphasize the multiplicative objective due to our motivation of avoiding catastrophe, but our subconstant additive regret bound may also be of interest. In summary, the combination of a mentor and local generalization allows us to reduce the regret by an entire factor of $T$, resulting in subconstant regret (multiplicative or additive) instead of the typical sublinear regret.
@@ 2 Related work
Despite the ubiquity of irreparable/irreversible costs in the real world, theoretical work on this topic remains limited. This may be due to the fundamental modeling question of how the agent should learn about novel inputs or actions without actually trying them.
The most common approach is to allow the agent to ask for help. This alone is insufficient, however: the agent must have some way to decide when to ask for help. A popular solution is to perform Bayesian inference on the world model, but this has two tricky requirements: (1) a prior distribution which contains the true world model (or an approximation), and (2) an environment where computing (or approximating) the posterior is tractable. A finite set of possible environments satisfies both conditions but is unrealistic in many real-world scenarios. In contrast, our algorithm can handle an uncountable policy class and a continuous unbounded input space, which is crucial for many real-world scenarios in which one never sees the exact same input twice.
Bayesian inference combined with asking for help is studied by Cohen et al. (2021); Cohen & Hutter (2020); Kosoy (2019); Mindermann et al. (2018). It is also worth mentioning Hadfield-Menell et al. (2017); Moldovan & Abbeel (2012); Turchetta et al. (2016), who utilize Bayesian inference in the context of safe (online) reinforcement learning without asking for help (and without regret bounds).
We are only aware of two papers which theoretically address irreversibility without Bayesian inference: Grinsztajn et al. (2021) and Maillard et al. (2019). The former proposes to sample trajectories and learn reversibility based on temporal consistency between states: intuitively, if $s_{1}$ always precedes $s_{2}$, we can infer that $s_{1}$ is unreachable from $s_{2}$. Although the paper theoretically grounds this intuition, there is no formal regret guarantee. The latter presents an algorithm which asks for help in the form of rollouts from the current state. However, the regret bound and number of rollouts are both linear in the worst case, due to the dependence on the $\gamma^{*}$ parameter which roughly captures how bad an irreversible action can be. In contrast, our algorithm achieves good regret even when actions are maximally bad.
To our knowledge, we are the first to provide an algorithm which formally guarantees avoidance of catastrophe (with high probability) without Bayesian inference. We are also not aware of prior results comparable to our negative result, including in the Bayesian regime.
Constrained Markov Decision Processes (CMDPs).
CMDPs (Altman, 2021) require the agent to maximize reward while also satisfying safety constraints. See Gu et al. (2023); Zhao et al. (2023); Wachi et al. (2024) for surveys. The two most relevant papers are Liu et al. (2021) and Stradi et al. (2024), both of which provide algorithms guaranteed to satisfy initially unknown safety constraints. Since neither paper allows external help, they require strong assumptions to make the problem tractable: the aforementioned results assume that the agent (1) knows a strictly safe policy upfront (i.e., a policy which satisfies the safety constraints with slack), (2) is periodically reset, and (3) observes the safety costs. In contrast, our agent has no prior knowledge, is never reset, and never observes payoffs.
@@
See Cesa-Bianchi & Lugosi (2006) and Chapter 21 of Shalev-Shwartz & Ben-David (2014) for introductions to online learning. A classical result states that sublinear regret is possible iff the hypothesis class has finite Littlestone dimension (Littlestone, 1988). However, even some simple hypothesis classes have infinite Littlestone dimension, such as the class of thresholds on $[0,1]$ (Example 21.4 in Shalev-Shwartz & Ben-David (2014)). Recently, Haghtalab et al. (2024) showed that if the adversary only chooses a distribution over inputs rather than the precise input, only finite VC dimension (Vapnik & Chervonenkis, 1971) is needed for sublinear regret. Specifically, they assume that each input is sampled from a distribution whose concentration is upper bounded by $\frac{1}{\sigma}$ times the uniform distribution. This framework is known as smoothed analysis, originally due to Spielman & Teng (2004).
@@
Although online learning traditionally studies the sum of payoffs, there is some work on maximizing the product of payoffs, or equivalently the sum of logarithms (Chapter 9 of Cesa-Bianchi et al. (2017)). However, these regret bounds are still sublinear in $T$, in comparison to our subconstant regret bounds. Also, that work still assumes that payoffs are observed on every time step, while our agent only receives feedback in response to queries (Figure 1).
Barman et al. (2023) recently provided regret bounds for a multiplicative objective in a multi-armed bandit problem, but their objective is the geometric mean of payoffs instead of the product. Interpreted in our context, their regret bounds imply that the average chance of catastrophe goes to zero, while we guarantee that the total chance of catastrophe goes to zero. This distinction is closely related to the difference between subconstant and sublinear regret discussed in Section 1.3.
Active learning and imitation learning.
Our assumption that the agent only receives feedback in response to queries falls under the umbrella of active learning (Hanneke, 2014). This contrasts with passive learning, where the agent receives feedback automatically. Although ideas from active learning could be useful in our domain, we are not aware of any results from that literature which account for irreversible costs. The process of the agent learning from a mentor is also reminiscent of imitation learning (Osa et al., 2018), but we are not aware of any relevant technical implications.
@@ 3 Model
Inputs.
Let $\mathbb{N}$ refer to the set of strictly positive integers and let $T\in\mathbb{N}$ be the time horizon. Let $\boldsymbol{x}=(x_{1},x_{2},\dots x_{T})\in\mathcal{X}^{T}$ be the sequence of inputs. In the adversarial setting, each $x_{t}$ can have arbitrary dependence on the events of prior time steps. In the smoothed setting, the adversary only chooses the distribution $x_{t}$ from which $x_{t}$ is sampled. Formally, a distribution $\mathcal{D}$ over $\mathcal{X}$ is $\sigma$-smooth if for any $X\subseteq\mathcal{X}$, $\mathcal{D}(X)\leq\frac{1}{\sigma}U(X)$. (In the smoothed setting, we assume that $\mathcal{X}$ supports a uniform distribution $U$ [^6] For example, $\mathcal{X}$ having finite Lebesgue measure is sufficient. Note that this does not imply boundedness. Alternatively, $\sigma$-smoothness can be defined with respect to a different distribution, as long as the Radon-Nikodym derivative is uniformly bounded; see Definition 1 of Block et al. (2022).) If each $x_{t}$ is sampled from a $\sigma$-smooth $\mathcal{D}_{t}$, we say that $\boldsymbol{x}$ is $\sigma$-smooth. The sequence $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ can still be adaptive, i.e., the choice of $\mathcal{D}_{t}$ can depend on the events of prior time steps.
Actions and payoffs.
Let $\mathcal{Y}$ be a finite set of actions. There also exists a special action $\tilde{y}$ which corresponds to querying the mentor. For $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$. On each time step $t\in[T]$, the agent must select action $y_{t}\in\ \mathcal{Y}\cup\{\tilde{y}\}$, which generates a payoff. Let $\boldsymbol{y}=(y_{1},\dots,y_{T})$. We allow the payoff function to vary between time steps: let $\boldsymbol{\mu}=(\mu_{1},\dots,\mu_{T})\in(\mathcal{X}\times\mathcal{Y}\to[0,1])^{T}$ be the sequence of payoff functions. [^7] Variation in $\mu_{t}$ is not a core part of our problem, but we allow it because the exact same proof holds. Since the mentor policy is fixed, local generalization limits how much $\mu_{t}$ can vary between time steps. Then $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the agent’s payoff at time $t$. Like $\boldsymbol{\mathcal{D}}$, we allow $\boldsymbol{\mu}$ to be adaptive. Unless otherwise noted, all expectations are over the agent’s randomization (if any) and the randomization in $\boldsymbol{x}$ (if any).
@@
The mentor is endowed with a (not necessarily optimal) policy $\pi^{m}:\mathcal{X}\to\mathcal{Y}$. When action $\tilde{y}$ is chosen, the mentor informs the agent of the action $\pi^{m}(x_{t})$ and the agent obtains payoff $\mu_{t}(x_{t},\pi^{m}(x_{t}))$. For brevity, let $\mu_{t}^{m}(x)=\mu_{t}(x,\pi^{m}(x))$. The agent never observes payoffs: the only way to learn about $\boldsymbol{\mu}$ is by querying the mentor.
We would like an algorithm which becomes “self-sufficient” over time: the rate of querying the mentor should go to 0 as $T\to\infty$, or equivalently, the cumulative number of queries should be sublinear in $T$. Formally, let $Q_{T}(\boldsymbol{\mu},\pi^{m})=\{t\in[T]:y_{t}=\tilde{y}\}$ be the random variable denoting the set of time steps with queries. Then we say that the (expected) number of queries is sublinear in $T$ if $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}(\boldsymbol{\mu},\pi^{m})|]\in o(T)$. In other words, there must exist $g:\mathbb{N}\to\mathbb{N}$ such that $g(T)\in o(T)$ and $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}(\boldsymbol{\mu},\pi^{m})|]\leq g(T)$. [^8] One could instead consider the worst-case number of queries, but this distinction does not affect whether subconstant regret is achievable (Proposition E.2). For brevity, we will usually write $Q_{T}=Q_{T}(\boldsymbol{\mu},\pi^{m})$.
@@
We assume that the mentor policy permits “local generalization”. Informally, if the agent is given an input $x$, taking the mentor action for a similar input $x^{\prime}$ is almost as good as taking the mentor action for $x$. Formally, we assume $\mathcal{X}\subseteq\mathbb{R}^{n}$ and there exists $L>0$ such that for all $x,x^{\prime}\in\mathcal{X}$ and $t\in[T]$, $|\mu_{t}^{m}(x)-\mu_{t}(x,\pi^{m}(x^{\prime}))|\leq L||x-x^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance. [^9] One could also allow a generic metric space; our assumption of $\mathcal{X}\subseteq\mathbb{R}^{n}$ is only for convenience. This represents the ability to transfer knowledge between similar inputs:
$|\underbrace{\mu_{t}(x,\pi^{m}(x))}_{\text{Taking the right action}}-\underbrace{\mu_{t}(x,\pi^{m}(x^{\prime}))}_{\text{Using what you learned in $x^{\prime}$}}|\ \ \ \leq\underbrace{L||x-x^{\prime}||}_{\text{Similarity between $x$ and $x^{\prime}$}}$
Borrowing knowledge from similar experiences seems fundamental to learning and is well-understood in the psychology literature (Esser et al., 2023) and education literature (Hajian, 2019).
Crucially, our input space can be any feature embedding of the agent’s situation, not just its physical positioning. Our algorithms will not require knowledge of the feature embedding and do not need to know $L$, so it suffices that there exists some feature embedding which satisfies local generalization. The agent does not even need to know which embedding it is. Finally, local generalization implies the more familiar Lipschitz continuity for an optimal mentor (Proposition E.1).